OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 3x - 3y = 4. slope = -1; y intercept at (0, -four thirds) slope = 1; y intercept at (0, -four thirds) slope = 3; y intercept at (0, -1) slope = -3; y intercept at (0, 1)

OpenStudy (anonymous):

you need to put it in y = mx + b form.....then m is the slope and b is the y intercept

OpenStudy (anonymous):

That's what I get confused with lol

OpenStudy (anonymous):

3x - 3y = 4 -3y = -3x + 4 (multiply by -1 to get 3y positive) 3y = 3x - 4 are you with me so far ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3y = 3x - 4 (now divide by 3) y = 3/3x - 4/3 y = x - 4/3 so your answer is : slope 1, and y intercept (0,-4/3)

OpenStudy (anonymous):

Thank you Kellie may you help me in another one??

OpenStudy (anonymous):

possibly...what you got

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 6) and parallel to y = one halfx - 8. y = one halfx + 7 y = one halfx + 5 y = 2x + 7 y = 2x + 5

OpenStudy (anonymous):

if you are looking for a parallel line, the slope will be the same as the original equation. In this case, the slope is 1/2. Understand ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now we will use y = mx + b.....m = 1/2 and we can use points (-2,6) now lets sub... 6 = 1/2(-2) + b can you solve for b ?

OpenStudy (anonymous):

No .-.

OpenStudy (anonymous):

6 = 1/2(-2) + b 6 = -1 + b 6 + 1 = b 7 = b therefore, your parallel equation is : y = 1/2x + 7

OpenStudy (anonymous):

Thanks (: I get it now .

OpenStudy (anonymous):

good deal :) glad to help